OpenStudy (kitkat1):

Medal for best answer!!!!!! Consider the functions f(x) = 3x − 2 and g(x) = x2 − 1. What is the value of f[g(5)]? 312 70 168 −1

rebeccaxhawaii (rebeccaxhawaii):

plug in 5 into the g(x) fomula

rebeccaxhawaii (rebeccaxhawaii):

so \[5^{2}-1\] simplify 25-1 giving you 24

rebeccaxhawaii (rebeccaxhawaii):

so then f(24) because x equals g(x)

rebeccaxhawaii (rebeccaxhawaii):

so f(24)=3(24)-2 72-2 70 is your final answer

OpenStudy (kitkat1):

Are you sure?

rebeccaxhawaii (rebeccaxhawaii):

are u sure you're gonna ask me that ?

OpenStudy (kitkat1):

Okay fine can you help me with last one?

OpenStudy (kitkat1):

Which expression is equivalent to the area of metal sheet required to make this square-shaped traffic sign? (5 points) A square shaped traffic sign is shown with length of one side labeled as 2x minus 1. 4x2 − 1 4x2 + 1 4x2 + 4x − 1 4x2 − 4x + 1

rebeccaxhawaii (rebeccaxhawaii):

pls medal and make a new question .. tag me then i shall help you

OpenStudy (kitkat1):

I cant open a new post becuase my computer is lagging I can only work on this post

OpenStudy (anonymous):

Okay, x=(-2) f(-2) =3*(-2) - 2 f(-2) = -6 -2 f(-2) = -8 g(-2) = (-2)^2 − 1 g(-2) = 4 -1 g(-2) = 3 (f-g) = -8 - 3

rebeccaxhawaii (rebeccaxhawaii):

area formula is L x W so (2x-1)(2x-1) then use the distributive property